OpenStudy (anonymous):

sin theta = 0.5 solve 0≤theta≤360

OpenStudy (mathmale):

So: you're told that the sine of the angle "theta" is 0.5, or 1/2. First: In which quadrants is the sine positive?

OpenStudy (mathmale):

Think of the 30-60-90 triangle and the related lengths of its sides. Could you use this triangle to determine which 2 angles have the sine 0.5?

OpenStudy (anonymous):

\[sin(\theta)=0.5\] to this quadrant we have 2 possible values. \[\theta=arcsin(0.5)\] \[\boxed{\boxed{\theta_1=30^o~~~~or~~~~\theta_2=150^o}}\]